OpenStudy (anonymous):

Equation #1 6 - x = 5x + 30 Equation #2 4x - 8 = 3x + 2 The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. None of the statements above describe solutions to equations shown.

OpenStudy (anonymous):

do u know it

OpenStudy (anonymous):

is it c

OpenStudy (radar):

Solve each equation and compare your answer. 6-x=5x+30 -x -5x = 30 - 6 -6x = 24 x= -4 Solve the second one and compare the results.

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the inequality y > -x + 5? (1, -10) (0, 5) (5, -1) (1, 9)

OpenStudy (anonymous):

its b